OpenStudy (anonymous):

if f(x) =3x + 2, find f(4)

OpenStudy (anonymous):

f(4) = 3*4 + 2 = 12 + 2 = 14

OpenStudy (anonymous):

u got it

OpenStudy (anonymous):

yes, Thank you so much! (: @jose12345

OpenStudy (anonymous):

no problem i need to go talk to u tomorrow at what time i will be alone message me on email if u want take care

OpenStudy (anonymous):

sure i will help u want u need

OpenStudy (anonymous):

u there